Find the indicated limit if exists. lim x2-16 3) x 4 x 4 A) The limit fails to exist. 08 B) 4 DO
Answer #1

solve: Criveni lim 2²16 X-4 X-4 2 lim 2 - 4 Piid+b= 6+b)la-b? So; lim (2-4) (x+4) (x-4) a lim (x+4) x + 4 Now, plug in x=-4 S